OpenStudy (anonymous):

Two cars leave town at the same time traveling in opposite directions. One car travels at a rate of 60 miles per hour. The other car travels at a rate of 70 miles per hour. How long will it take them to arrive at a point where they are 650 miles apart?

OpenStudy (anonymous):

@helper99

OpenStudy (anonymous):

D=RT 650=(60+70)T 650=130T T=650/130 T=5 HOURS ThEY WILL BE 650 MILES APART. PROOF: 650=(60+70)5 650=130*5 650=650

OpenStudy (ciarán95):

There's one key formula we need here I think: DISTANCE = SPEED x TIME So, we want to know how long it will be until the distance covered both cars, when added together, will be 650 miles. We know that both cars will be travelling for the same length of time, but just at different speeds. Let's call the unknown length of time x. So, we'll say Car A is travelling at 60 miles per hour and that car B is travelling at 70 miles per hour. We know that, at time 'X' (DISTANCE TRAVELED BY CAR A) + (DISTANCE TRAVELED BY CAR B) = 650 MILES or, ((SPEED OF CAR A)x(TIME CAR A HAS BEEN MOVING)) + ((SPEED OF CAR B)x(TIME CAR B HAS BEEN MOVING)) = 650 MILES So, we know that TIME CAR A HAS BEEN MOVING = TIME CAR B HAS BEEN MOVING = X And that SPEED OF CAR A = 60 m/h and SPEED OF CAR B = 70 m/h. So, we substitute in these values into the equation above: (60 miles/hour)(X) + (70 miles/hour)(X) = 650 miles We can add the two terms on the left-hand side of the equation: (130 miles/hour)(X) = 650 miles If we divide both sides of the equation by 130 miles/hour, the 130 on top and below on the left-hand side will cancel: \[\frac{ 130m/h }{130m/h}X = \frac{ 650 m }{ 130m/h }\] \[X = \frac{ 650 m }{ 130m/h }\] which, when you work it out, will give you the length of time, IN HOURS, that the two cars must travel to be 650 miles apart. :)